How long will it take 1.5 Ampere of current to liberate a layer of Coulomb 0.1mm thickness on both sides 12 by 12cm of the electrode, if the electrochemical equivalent is 0.003cm by 9.0 density.


1
Expert's answer
2021-12-02T10:08:13-0500

The equation is


"t=\frac{M}{IZ}=\frac{\rho V}{IZ},\\\space\\
t=\frac{9000·(0.1·10^{-3}·0.12^2)}{1.5·0.00003}=288\text{ s}."